# Extreme Forward Scattering Observed in Disk-Averaged Near-Infrared Phase Curves of Titan

### v1

@@ 1 Introduction
Observations of exoplanets have revealed an impressive diversity in planetary atmospheres. The atmospheres of many Jupiter- to Neptune-sized exoplanets have been successfully studied by looking at features in transit and/or thermal emission spectra (e.g., Madhusudhan et al., 2020; Kesseli et al., 2022; Zhang et al., 2022; Madhusudhan et al., 2023; Rustamkulov et al., 2023). For a subset of worlds — especially hot Jupiters — observations of their changing thermal brightness with star-planet-observer (i.e., phase) angle has helped to constrain key aspects of atmospheric thermal structure and circulation (for a convenient summary of targets, see Table 1 in Parmentier & Crossfield, 2018).
In reflected-light direct imaging, current telescopes are generally unable resolve planetary companions and study phase curves, especially for rocky exoplanets (see, e.g., Wang et al., 2017). In the future, though, NASA’s under-development Habitable Worlds Observatory (HWO; Feinberg et al., 2024) will provide the high-contrast imaging capabilities required to study a wide range of planet types, including clement terrestrial worlds. Critically, repeat visits to planetary systems and targets will enable HWO to build up phase curves for many worlds over the duration of the mission.
Studying reflected-light phase curves can constrain surface and/or atmospheric properties of a planet. In general, an object’s brightness will decrease as phase angle increases, as progressively less of the planetary disk is illuminated from the observer’s perspective. Moving beyond simple illumination effects, the shape of phase curves can be affected by phenomena that can reveal details of the atmospheric and/or surface state. For example, phase curves of Venus contain a feature at high phase angles due to sulfuric acid droplets in the atmosphere (Arking & Potter, 1968), while optical phase curves of Titan show that the moon’s haze is responsible for forward scattering of optical light (García Muñoz et al., 2017). Phase curves of solar system objects can further serve as testing grounds for future direct observations of exoplanets and the production of their phase curves, as has been done with both Jupiter (Mayorga et al., 2016; Heng & Li, 2021) and Saturn (Hedman & Stark, 2015; Dyudina et al., 2016).
@@
Most fundamentally, an ocean glint signature in a planetary phase curve is revealed through forward scattering. Thus, the previously mentioned haze forward scattering detections from García Muñoz et al. (2017) could present a potential false positive for glint. This connection between glint and Titan is particularly interesting given the collection of studies that use glint from Titan to study its seas (Stephan et al., 2010; Barnes et al., 2011; Soderblom et al., 2012; Barnes et al., 2013, 2014). The work presented here further explores aerosol forward scattering in Titan phase curves.
In the solar system, Titan stands out among planetary objects as one of only a handful of rocky/icy bodies with thick atmospheres. Due to the large semi-major axis of the Saturn system in comparison to that of the Earth, both ground and Earth-orbiting telescopes can only view Titan at phase angles below about . Fortunately, access to viewing angles not obtainable from Earth was provided by NASA’s Cassini spacecraft, which made orbital observations of Saturn, its rings, and its moons from 2004 to 2017 (Titan Discipline Working Group, 2019). Cassini made regular flybys of Titan, taking data of the moon in the ultraviolet, visible, infrared, and radio wavelength regimes. Among the instruments on board the probe was the Visual and Infrared Mapping Spectrometer (VIMS) instrument, which took spatially resolved images of Titan in the optical (0.3–0.9 m) and near-infrared (0.9–5.1 m) wavelength regimes (Brown et al., 2004).
The primary focus of this work is to use Cassini/VIMS observations of Titan to generate phase curves that span the near-infrared wavelength regime, thereby complementing the optical studies of García Muñoz et al. (2017). Section 2 covers data acquisition and analysis using available resources, as well as the pipeline we produced to reduce spatially resolved images to disk-averaged reflectivity measures. In Section 3, near-infrared reflected-light phase curves of Titan are presented and key features are noted. Section 4 explores possible explanations of the two phenomena mentioned in Section 3, and compares our findings to those of García Muñoz et al. (2017). Finally, Section 5 will summarize the findings of this work, as well as connect these findings to ongoing research of exoplanets.
@@ 2.1 File Vetting and Calibration
Total invalid cubes | 57,849 |
@@
Data were acquired using the PDS Image Atlas on NASA’s Planetary Data System. In the rest of this paper, downloaded files are referred to as “cubes.” Cubes contain general information about the time and duration of the observation, as well as the location and orientation of Cassini at time of acquisition. Each cube includes 352 two-dimensional arrays—one array for each VIMS band—and each array contains spatially-resolved counts accumulated during the cube’s exposure time. Arrays 1–96 correspond to VIMS bands in optical light, while arrays 97–352 correspond to bands at infrared (IR) wavelengths. Our analyzed cubes were acquired throughout the full duration of the Cassini mission (2004–2017), were taken at phase angles from 2°to 165°, and contained imaging data at all infrared bands covered by the VIMS-IR camera (0.9–5.1 m). A total of 62,376 cubes of Titan were downloaded.
@@
where radians is the field of view of a VIMS pixel on a side, is the radius of Titan in km, and is the distance of the spacecraft at time of observation. By letting and solving for , we get a lower distance limit of approximately 180,000 km. Cubes missing data from more than 10% of VIMS bands were also removed. We then used a simple Canny edge detection method to test for structures in cube data, and excluded cubes showing no structure. Edge detection was also used to remove cubes not containing the full disk of Titan. Finally, remaining cubes were manually inspected to ensure they met our criteria. Table 1 breaks down the different criteria used to sort the cubes, in the order they were applied, and the amount of cubes removed by each criterion. Once the filtering process was finished, our dataset contained 4,492 cubes.
All cubes were passed through the multi-step calibration process described by Le Mouélic et al. (2019) using the USGS ISIS3 software package. The calibration process formats each cube to be ISIS3-compliant, performs background subtraction and noise filtering, removes sources of error such as imaging artifacts and cosmic rays, and adds SPICE data used for calculating spacecraft position and pointing data. The calibration also converted data numbers, which are correlated with photon counts, to spatially-resolved values, a unit of measurement of an object’s reflectivity (see Section 2.3). Cubes were calibrated this way in order to be compatible with the pyvims Python library, the primary tool used for the analysis of cube data. The pyvims software suite (described in Le Mouélic et al., 2019) facilitates the analysis of VIMS cubes and can extract from a cube the full suite of ephemeris and pointing quantities, including sub-spacecraft coordinates and phase angle.
VIMS data taken in the 0.9–1.3 m wavelength range suffered from extensive saturation issues, especially at high phase angles due to scattering effects. The cause is likely longer integration times – cubes with longer exposure times showed a greater degree of saturation. The phase curves produced from these data show greater spread and do not closely resemble a continuous curve. As a result, we are unable to properly analyze most of these these data. These issues extended to a few cubes at longer wavelengths at higher phase angles, and these cubes were excluded from our dataset via manual sorting.
@@ 2.2 Automated Disk Detection
The automatic detection of Titan within images was achieved using a pipeline written in Python, and mirrors the disk detection process used by Strauss et al. (2024). A visual summary of the process can be seen in Figure 1. First, the radius of Titan is determined by adding the wavelength-dependent atmospheric height of Titan (Robinson et al., 2014) to the solid body radius of km (Zebker et al., 2009). Effective height data at wavelengths corresponding to VIMS-IR bands 345–352 were absent, so the effective height used for band 344 was applied to these bands as well. Then Titan’s radius in units of VIMS pixels is calculated as , where is the average surface resolution of the image in kilometers per pixel as determined with pyvims. In cases where the phase angle of the image is above , the atmosphere appears much brighter due to forward scattering of light by atmospheric aerosols, which increases the apparent radius of Titan, so is increased by a pixel to account for this effect. Increasing the radius by more than 1 pixel failed to improve sensitivity to increased atmospheric brightness.
Next, Canny edge detection (Canny, 1986) is applied to the image. Canny edge detection algorithms are a family of algorithms that identify edges and boundaries in pixelated images. In the case of images, Canny edge detection returns an array that identifies pixels which likely lie on the edge of the disk of Titan.
@@
A circle Hough transform is then used to locate the (approximate) center of the disk (Xie & Ji, 2002). In the circle Hough transform, an accumulator array with the same dimensions as the image is created. Circles with a radius of pixels are overlaid on the array, centered on each edge pixel location in the edge pixel array, with each pixel covered by a circle having its accumulator array value incremented. After each circle has been overlaid, the pixel with the highest value is taken to be the center of the disk, and an array is created where each pixel within pixels of the center pixel identified above has a value of 1, and all other pixels have a value of 0. By multiplying the original image array by this masking array, the resulting array only contains data from pixels lying on the disk of Titan.
2.3 I/F Calculation
@@
, also referred to as the radiance factor, is a dimensionless quantity in planetary science that measures the ability of a celestial object to reflect incident radiation. The cube calibration process described above converts photon counts from the raw Cassini data to values. Because we aim to simulate point-like observations of Titan rather than the spatially resolved observations provided by Cassini, we convert the spatial values across Titan’s disk to a single “disk averaged” quantity. The disk-averaged of the disk of Titan is calculated as,
@@
where is the distance from Cassini to the center of Titan in km, is the solid angle of a VIMS pixel in steradians, is the apparent radius of Titan in km, and the sum is over individual values of all pixels on the disk.
@@ 3 Results
Figure 4 shows the disk-averaged of Titan as a function of wavelength and phase angle. These disk-averaged values are not normalized to full phase, so are not formal phase functions. Cubes were binned by phase angle into one-degree bins, and for bins with multiple cubes an average I/F value was computed. Vertical black stripes represent phase angle bins with no viable cubes. The region on the right edge of the figure indicates where significant forward scattering causes Titan to appear bright, as discussed above. These regions largely correspond to continuum outside of Titan’s atmospheric absorption features.
Figure 5 shows a selection of phase curves at select wavelengths produced by our pipeline. The phase curve of Titan is decidedly not Lambertian, in contrast to those of other major Saturnian satellites which are better-approximated by a Lambert phase function (Buratti & Veverka, 1984). At every wavelength, the brightness of Titan follows a smooth curve that initially decreases as the phase angle increases from roughly to . At phase angles above about , the brightness of the disk then increases due to forward scattering of sunlight by the atmosphere. At phase angles above about , Titan appears brighter than when observed at near-full phase. While this effect has been observed in the 0.3–1 m range before (Tomasko et al., 2008; Doose et al., 2016; García Muñoz et al., 2017), our analysis confirms that significant forward scattering occurs at most infrared VIMS bands. The larger phase curve spread at smaller phase angles in the 2 m plot in Figure 5 is apparent in all phase curves recorded at wavelengths with sensitivity to the deep atmosphere and/or surface and will be discussed later.
Polynomial fits of degree 10 to select phase curves at wavelengths between 0.88m and 5.12m are shown in Figure 6, normalized to lowest-phase I/F values. All phase curves show a gradual decrease from roughly 2–80°before leveling out. Significant brightness surges occur beginning around 130°. The effect is weakest at shorter wavelengths, though for some bands this muting may be due to aforementioned saturation issues. Some normalized phase curves taken at wavelengths between 3 and 4m show nearly an order of magnitude increase between near-full and crescent phases as a result of significant forward scattering by atmospheric aerosols.
Figures 7 and 8 emphasize the value of disk-averaged Titan observations as an analog for a hazy exoplanet via comparisons to an Earth phase curve and Earth color-color data, respectively. Published measurements of Earth’s phase curve at wavelengths corresponding to VIMS do not exist, so the broadband visible (0.4–0.7 m) Earthshine data are shown (Qiu et al., 2003; Pallé et al., 2003). Similarly, phase-dependent color measurements for Earth at VIMS-analogous wavelengths do not exist, so we adopt high-fidelity, phase-dependent simulations from the Virtual Planetary Laboratory 3-D Spectral Earth Model (Robinson et al., 2010).
@@ 4.1 Phase Curve Structure and Comparisons to Optical Results
At all wavelengths, the phase curves presented here share the same general structure. Near full-phase, disk-averaged starts modest and decreases as phase angle increases, with minimum disk-averaged measurements occurring near °. Curves then sharply increase, beginning around 140° phase. Maximum disk-averaged measurements occur at high phase at nearly all wavelengths; at the shortest wavelengths accessible to VIMS, disk-averaged at lowest and highest phase are comparable. At atmospheric window wavelengths (e.g., the 2 m plot in Figure 5), enhanced backscattering due to the fewer scattering events causes stronger brightening between quadrature phase and full phase.
@@
The wavelength dependence of the normalized crescent phase peak is primarily controlled by aerosol scattering and gas absorption optical depths. Previous estimates on the size of Titan aerosols are on the order of 1 m (Rages et al., 1983; Waite et al., 2007; Tomasko et al., 2008), so that scattering from these aerosols at near-infrared wavelengths is in the Mie regime with strong forward-scattering peaks and decreasing optical depth with increasing wavelength (Hansen & Travis, 1974). The strong forward peak in the aerosol single-scattering phase function explains the general behavior of larger disk-averaged values at crescent phases versus near full phase given the ubiquity of haze aerosols in Titan’s atmosphere (García Muñoz et al., 2017). Increasing atmospheric aerosol transparency at longer wavelengths causes scattering to tend towards the single-scattering regime, which explains the general trend with wavelength in Figure 6. The figure breaks from this trend around 4.3 m, potentially due to local minima in the absorptivity of tholins around these wavelengths (see Figure 4 of Khare et al., 1984). At wavelengths with strong gas absorption, photons incident at the near-full phase geometry would typically require multiple scatterings to escape the atmosphere so are, instead, absorbed along such a path. At crescent phases, though, only a limited number of scattering events are required to direct photons towards the observer (spacecraft), making observations at these wavelengths and phase less sensitive to gas absorption in the deeper atmosphere and more sensitive to single-scattered radiation. A comprehensive spectral model of aerosol scattering could further explain these results but falls beyond the scope of this work.
@@
The general shape of near-infrared phase curves produced with our pipeline matches that of the optical phase curves produced by García Muñoz et al. (2017). A comparison of our results to data from this earlier work is shown in Figure 9. Both sets demonstrate intense forward scattering from Titan’s atmospheric aerosols. The onset of brightness surges occur at similar phase angles in our phase curves (135°) as in those of García Muñoz et al. (2017) (150°). However, the relative strength of forward scattering is evidently much different between the two wavelength regimes. Several near-infrared phase curve fits across our sampled phase angle range (2–165°) show high- to low-phase disk-averaged ratios of 10 or more; optical light phase curve ratios across the same intervals are of order unity (consistent with results at the shortest VIMS-IR wavelengths). The differences in the extent of the crescent phase peak in Titan’s phase curves at optical versus near-infrared wavelengths aligns with the earlier physical explanation rooted in haze scattering.

### v3

@@ 1 Introduction
Observations of exoplanets have revealed an impressive diversity in planetary atmospheres. The atmospheres of many Jupiter- to Neptune-sized exoplanets have been successfully studied by looking at features in transit (e.g., Madhusudhan et al., 2020; Kesseli et al., 2022; Zhang et al., 2022; Madhusudhan et al., 2023; Rustamkulov et al., 2023), thermal emission (Charbonneau et al., 2005; Knutson et al., 2008; Kreidberg et al., 2014; Stevenson et al., 2014), and/or reflected light (Esteves et al., 2015; Hoeijmakers et al., 2018; Lendl et al., 2020; Hooton et al., 2022; Brandeker et al., 2022) spectra. For a subset of worlds — especially hot Jupiters — observations of their changing thermal brightness with star-planet-observer (i.e., phase) angle has helped to constrain key aspects of atmospheric thermal structure and circulation (for a convenient summary of targets, see Table 1 in Parmentier & Crossfield, 2018).
In reflected-light direct imaging, current telescopes are generally unable to resolve planetary companions and study phase curves, especially for rocky exoplanets (see, e.g., Wang et al., 2017). In the future, though, NASA’s under-development Habitable Worlds Observatory (HWO; Feinberg et al., 2024) will provide the high-contrast imaging capabilities required to study a wide range of planet types, including clement terrestrial worlds. Critically, repeat visits to planetary systems and targets will enable HWO to build up phase curves for many worlds over the duration of the mission.
Studying reflected-light phase curves can constrain surface and/or atmospheric properties of a planet. In general, an object’s brightness will decrease as phase angle increases, as progressively less of the planetary disk is illuminated from the observer’s perspective. Moving beyond simple illumination effects, the shape of phase curves can be affected by phenomena that can reveal details of the atmospheric and/or surface state. For example, phase curves of Venus contain a feature at high phase angles due to sulfuric acid droplets in the atmosphere (Arking & Potter, 1968), while optical phase curves of Titan show that the moon’s haze is responsible for forward scattering of optical light (Garc´ıa Muñoz et al., 2017). Phase curves of solar system objects can further serve as testing grounds for future direct observations of exoplanets and the production of their phase curves, as has been done with both Jupiter (Mayorga et al., 2016; Heng & Li, 2021) and Saturn (Hedman & Stark, 2015; Dyudina et al., 2016).
@@
Most fundamentally, an ocean glint signature in a planetary phase curve is revealed through forward scattering. Thus, the previously mentioned haze forward scattering detections from Garc´ıa Muñoz et al. (2017) could present a potential false positive for glint. This connection between glint and Titan is particularly interesting given the collection of studies that use glint from Titan to study its seas (Stephan et al., 2010; Barnes et al., 2011; Soderblom et al., 2012; Barnes et al., 2013, 2014). The work presented here further explores aerosol forward scattering in Titan phase curves.
In the solar system, Titan stands out among planetary objects as an example of a rocky/icy body with a thick atmosphere. Due to the large semi-major axis of the Saturn system in comparison to that of the Earth, both ground and Earth-orbiting telescopes can only view Titan at phase angles below about . Fortunately, access to viewing angles not obtainable from Earth was provided by NASA’s Cassini spacecraft, which made orbital observations of Saturn, its rings, and its moons from 2004 to 2017 (Titan Discipline Working Group, 2019). Cassini made regular flybys of Titan, taking data of the moon in the ultraviolet, visible, infrared, and radio wavelength regimes. Among the instruments on board the probe was the Visual and Infrared Mapping Spectrometer (VIMS) instrument, which took spatially resolved images of Titan in the optical (0.3–0.9 m) and near-infrared (0.9–5.1 m) wavelength regimes (Brown et al., 2004).
The primary focus of this work is to use Cassini/VIMS observations of Titan to generate phase curves that span the near-infrared wavelength regime, thereby complementing the optical studies of Garc´ıa Muñoz et al. (2017). Section 2 covers data acquisition and analysis using available resources, as well as the pipeline we produced to reduce spatially resolved images to disk-averaged reflectivity measures. In Section 3, near-infrared reflected-light phase curves of Titan are presented and key features are noted. Section 4 explores possible explanations of the two phenomena mentioned in Section 3, and compares our findings to those of Garc´ıa Muñoz et al. (2017). Finally, Section 5 will summarize the findings of this work, as well as connect these findings to ongoing research of exoplanets.
@@ 2.1 File Vetting and Calibration
| Total invalid cubes | 57,884 |
@@
Data were acquired using the PDS Image Atlas on NASA’s Planetary Data System. In the rest of this paper, downloaded files are referred to as “cubes.” Cubes contain general information about the time and duration of the observation, as well as the location and orientation of Cassini at time of acquisition. Each cube includes 352 two-dimensional arrays—one array for each VIMS band—and each array contains spatially-resolved counts accumulated during the cube’s exposure time. Arrays 1–96 correspond to VIMS bands in optical light, while arrays 97–352 correspond to bands at infrared (IR) wavelengths (Brown et al., 2004). Our analyzed cubes were acquired throughout the full duration of the Cassini mission (2004–2017), were taken at phase angles from to , and contained imaging data at all infrared bands covered by the VIMS-IR camera (0.9–5.1 m). A total of 62,376 cubes of Titan were downloaded.
@@
where radians is the field of view of a VIMS pixel on a side, is the radius of Titan in km, and is the distance of the spacecraft at time of observation. By letting and solving for , we get a lower distance limit of approximately 180,000 km. Cubes missing data from more than 10% of VIMS bands were also removed.
We then used a simple Canny edge detection method to test for structures in cube data, and excluded cubes showing no structure. Canny edge detection algorithms are a family of algorithms that identify edges and boundaries in pixelated images (Canny, 1986). In the case of Titan images, Canny edge detection returns an array that identifies pixels which either lie on the edge of the disk of Titan or on the boundaries between geographical regions. By requiring that cubes have structure as determined by this algorithm, we exclude cubes that do not contain the disk of Titan, that zoom in on a subsection of the disk, or have erroneous data. Edge detection was also used to remove cubes not containing the full disk of Titan. Finally, remaining cubes were manually inspected to ensure they met our criteria. Table 1 breaks down the different criteria used to sort the cubes, in the order they were applied, and the amount of cubes removed by each criterion. Once the filtering process was finished, our dataset contained 4,492 cubes.
All cubes were passed through the multi-step calibration process described by Le Mouélic et al. (2019) using the USGS ISIS3 software package (Kelvin Rodriguez & Astrogeology Science Center, 2024). The calibration process formats each cube to be ISIS3-compliant, performs background subtraction and noise filtering, removes sources of error such as imaging artifacts and cosmic rays, and adds SPICE data. SPICE data includes information on the location and orientation of the spacecraft during its mission, as well as instrument details, target information and more (Acton, 1996; Acton et al., 2018). The calibration also converted data numbers, which are correlated with photon counts, to spatially-resolved values, a unit of measurement of an object’s reflectivity per steradian (see Section 2.3). Cubes were calibrated this way in order to be compatible with the pyvims Python library, the primary tool used for the analysis of cube data. The pyvims software suite (described in Le Mouélic et al., 2019) facilitates the analysis of VIMS cubes and can extract from a calibrated cube the full suite of ephemeris and pointing quantities, including sub-spacecraft coordinates and phase angle.
VIMS data taken in the 0.9–1.3 m wavelength range suffered from extensive saturation issues, especially at high phase angles. The cause is likely longer integration times – cubes with longer exposure times showed a greater degree of saturation. The phase curves produced from these data show greater spread and do not closely resemble a continuous curve. As a result, we are unable to properly analyze most of these data. We include these data in later figures for comparisons with previous works. Similar saturation issues extended to a few cubes at longer wavelengths at higher phase angles, and these cubes were excluded from our dataset via manual sorting.
@@ 2.2 Automated Disk Detection
The automatic detection of Titan within images was achieved using a pipeline written in Python, and mirrors the disk detection process used by Strauss et al. (2024). A visual summary of the process can be seen in Figure 1. First, the radius of Titan is determined by adding the wavelength-dependent atmospheric height of Titan (Robinson et al., 2014) to the solid body radius of km (Zebker et al., 2009). Effective height data at wavelengths corresponding to VIMS-IR bands 345–352 were absent, so the effective height used for band 344 was applied to these bands as well. Then Titan’s radius in units of VIMS pixels is calculated as , where is the average surface resolution of the image in kilometers per pixel as determined with pyvims. This radius is increased by 1 in cases where the phase angle of the image is above , because the atmosphere appears much brighter due to forward scattering of light by atmospheric aerosols, which increases the apparent radius of Titan. Increasing this radius by more than 1 pixel did not substantially change the disk-averaged measurements.
@@
Next, we used a Canny edge detection algorithm to identify pixels on the edge of Titan’s visible disk. A circle Hough transform is then used to locate the (approximate) center of the disk (Xie & Ji, 2002). In the circle Hough transform, an accumulator array with the same dimensions as the image is created. Circles with a radius of pixels are overlaid on the array, centered on each edge pixel location in the edge pixel array, with each pixel covered by a circle having its accumulator array value incremented. After each circle has been overlaid, the pixel with the highest value is taken to be the center of the disk, and an array is created where each pixel within pixels of the center pixel identified above has a value of 1, and all other pixels have a value of 0. By multiplying the original image array by this masking array, the resulting array only contains data from pixels lying on the disk of Titan.
2.3 Calculation
@@
The quantity is the product of an object’s geometric albedo, , and its planetary phase function, . In planetary science, is often used as a metric of a planetary object’s reflectivity, as it quantifies the ratio of the radiance received from an object to the irradiance that object receives. As a reflectance, it is a unitless quantity. The cube calibration process described above converts data numbers from the raw Cassini data to values, with units of inverse steradians. Because we aim to simulate point-like observations of Titan rather than the spatially resolved observations provided by Cassini, we convert the spatial values across Titan’s disk to a single “disk averaged” quantity. The disk-averaged for Titan is calculated as,
@@
where is the disk-averaged intensity-to-incident flux ratio, is the distance from Cassini to the center of Titan in km, is the solid angle of a VIMS pixel in steradians, is the solid body radius of Titan in km, and the sum is over individual values of all pixels on the disk. pyvims does not provide information on errors in individual pixel measurements, however integrating over the disk significantly reduces random errors.
@@ 3 Results
Figure 3 shows the value of for Titan as a function of wavelength and phase angle. These disk-averaged values are not normalized to full phase, so are not formal phase functions. Cubes were binned by phase angle into one-degree bins, and for bins with multiple cubes an average value was computed. Vertical black stripes represent phase angle bins with no viable cubes. The region on the right edge of the figure indicates where significant forward scattering causes Titan to appear bright, as discussed above. These regions largely correspond to continuum outside of Titan’s atmospheric absorption features.
Figure 4 shows a selection of phase curves at select wavelengths produced by our pipeline. The phase curve of Titan is decidedly not Lambertian, in contrast to those of other major Saturnian satellites which are better-approximated by a Lambert phase function (Buratti & Veverka, 1984). At every wavelength, the brightness of Titan follows a smooth curve that initially decreases as the phase angle increases from roughly to . At phase angles above about , the brightness of the disk then increases due to forward scattering of sunlight by the atmosphere. At phase angles above about , Titan appears brighter than when observed at near-full phase. While this effect has been observed in the 0.3–1 m range before (Tomasko et al., 2008; Doose et al., 2016; Garc´ıa Muñoz et al., 2017), our analysis confirms that significant forward scattering occurs at most infrared VIMS bands. The larger spread in at smaller phase angles in the 2 m plot in Figure 4 is apparent in all phase curves recorded at wavelengths with sensitivity to the deep atmosphere and/or surface and will be discussed later.
Polynomial fits of degree 10 to select phase curves at wavelengths between 0.88m and 5.12m are shown in Figure 5, normalized to lowest-phase values, thereby approximating the planetary phase function . All phase curves show a gradual decrease from roughly 2–80 before leveling out. Significant brightness surges occur beginning around 130°. The effect is weakest at shorter wavelengths, though for some bands this muting may be due to aforementioned saturation issues. Some normalized phase curves taken at wavelengths between 3 and 4m show nearly an order of magnitude increase between near-full and crescent phases as a result of significant forward scattering by atmospheric aerosols.
Figures 6 and 7 emphasize the value of disk-averaged Titan observations as an analog for a hazy exoplanet via comparisons to an Earth phase curve and Earth color-color data, respectively. Published measurements of Earth’s phase curve at wavelengths corresponding to VIMS do not exist, so the broadband visible (0.4–0.7 m) Earthshine data are shown (Qiu et al., 2003; Pallé et al., 2003). Similarly, phase-dependent color measurements for Earth at VIMS-analogous wavelengths do not exist, so we adopt high-fidelity, phase-dependent simulations from the Virtual Planetary Laboratory 3-D Spectral Earth Model (Robinson et al., 2010).
@@ 4.1 Phase Curve Structure and Comparisons to Optical Results
At all wavelengths, the phase curves presented here share the same general structure. Near full-phase, disk-averaged starts modest and decreases as phase angle increases, with minimum disk-averaged measurements occurring near °. Curves then sharply increase, beginning around 140° phase. Maximum disk-averaged measurements occur at high phase at nearly all wavelengths; at the shortest wavelengths accessible to VIMS, disk-averaged at lowest and highest phase are comparable, though this is partially due to aforementioned saturation problems at these wavelengths. At atmospheric window wavelengths (e.g., the 2 m plot in Figure 4), photons are scattered fewer times overall, in many cases only once, and the back-scattering peak in the haze’s scattering phase function (West & Smith, 1991) results in stronger brightening between quadrature phase and full phase.
@@
The wavelength dependence of the normalized crescent phase peak is primarily controlled by aerosol scattering and gas absorption optical depths. Previous estimates on the size of Titan aerosols are on the order of 1 m (Rages et al., 1983; Waite et al., 2007; Tomasko et al., 2008), and strong forward-scattering is to be expected when particle sizes are approximately the same size as the wavelength of incident light. However, a complete description of the scattering behavior of Titan’s hazes requires consideration of the fractal aggregate nature of these hazes (see West & Smith, 1991; Tomasko et al., 2008; Lavvas et al., 2010). The strong forward peak in the aerosol single-scattering phase function explains the general behavior of larger disk-averaged values at crescent phases versus near full phase given the ubiquity of haze aerosols in Titan’s atmosphere (Garc´ıa Muñoz et al., 2017). Increasing atmospheric aerosol transparency at longer wavelengths causes scattering to tend towards the single-scattering regime, which explains the general trend with wavelength in Figure 5. The figure breaks from this trend around 4.3 m, though the exact reason for this is unknown. At wavelengths with strong gas absorption, photons incident at the near-full phase geometry would typically require multiple scatterings to escape the atmosphere so are, instead, absorbed along such a path. At crescent phases, though, only a limited number of scattering events are required to direct photons towards the observer (spacecraft), making observations at these wavelengths and phase less sensitive to gas absorption in the deeper atmosphere and more sensitive to single-scattered radiation. A comprehensive spectral model of aerosol scattering could further explain these results but falls beyond the scope of this work.
@@
The general shape of near-infrared phase curves produced with our pipeline matches that of the optical phase curves produced by Garc´ıa Muñoz et al. (2017). A comparison of our results to data from this earlier work is shown in Figure 8. Both sets demonstrate intense forward scattering from Titan’s atmospheric aerosols. The onset of brightness surges occur at similar phase angles in our phase curves (135°) as in those of Garc´ıa Muñoz et al. (2017) (150°). However, the relative strength of forward scattering is evidently much different between the two wavelength regimes. Several near-infrared phase curve fits across our sampled phase angle range (2–165) show approximate values of 10 or more at high phase; optical light phase curve values across the same intervals are of order unity (consistent with results at the shortest VIMS-IR wavelengths). The differences in the extent of the crescent phase peak in Titan’s phase curves at optical versus near-infrared wavelengths aligns with the earlier physical explanation rooted in haze scattering.